# Night-Rider: Nocturnal Vision-aided Localization in Streetlight Maps Using Invariant Extended Kalman Filtering

### 01 Feb 2024 v1

@@ IV-A Implementation Details
The robot is equipped with a front-facing RGB camera ($30$ Hz $1280\times 720$ images), an odometer ($10$ Hz) and a Livox $360^{\circ}$ FOV LiDAR [^2] https://www.livoxtech.com/mid-360 ($10$ Hz) integrating a 6-axis IMU ($200$ Hz). The camera’s intrinsics and all extrinsics among sensors have been pre-calibrated. All experiments were performed on the Intel i7-1165G7 CPU. Note that only an IMU, an odometer, and a monocular camera are used for the proposed nocturnal vision-aided localization system.
@@
Figure 5: Trajectories of ground truth (gray dashed line), VINS-Odom (green) and Night-Rider (blue) in four real scenes.
@@ IV-B Evaluation of Localization
To thoroughly evaluate the performance of the proposed system in nocturnal environments, we collect data in four different real scenes with multiple distribution types of streetlights. In the first scene (1149.2 m) and the fourth scene (723.8 m), the streetlights distribute sparsely on both sides of the road. Partial streetlights lie on only one side of the road in the second scene (1070.9 m). The third scene (777.2 m) contains densely distributed streetlights. All streetlight maps of these scenes are constructed using the method described in Section III-A. The offline optimized and precisely aligned trajectories from the loop closure augmented FAST-LIO2 using the LiDAR on board are treated as ground truths and EVO [30] is used for evaluation. Note that there are very few open-source research works targeting nocturnal scenarios. Thus we first compare Night-Rider with the state-of-the-art framework, extended VINS-Mono [3] with the odometer[31] (named VINS-Odom). As shown in Fig. 5, Night-Rider can precisely estimate trajectories with extremely low drifts in all four scenes while VINS-Odom drifts over time and even fails to track in the third scene. The reason is that the unstable illumination in nighttime scenes makes it difficult to extract reliable feature points from highly bright or dark images. This instability results in drifts or tracking failures of VINS-Odom. Contrarily, Night-Rider is independent of feature points and exploits the stable streetlights for localization. Thus trajectories of Night-Rider align accurately with ground truths. Table. I depicts the evaluation results of absolute trajectory error (ATE) for translation and rotation parts. It shows that Night-Rider achieves remarkably accurate and robust localization in all scenes (<0.2% relative error of total trajectory length) and totally outperforms VINS-Odom, which further validates the effectiveness of our system.
@@
Figure 6: Trajectories in $x$, $y$ and $z$-axis of ground truth (gray dashed line), Night-Rider w/o Vision (red), Night-Rider w/o ME (green) and Night-Rider (blue) in four real scenes.
TABLE I: ATE for translation and rotation of VINS-Odom, Night-Rider w/o Vision, Night-Rider w/o ME, and Night-Rider in four scenes. The best is shown in bold.
@@ IV-C Tracking Lost and Degeneration Cases
Fig. 7 shows that our system can successfully recover pose estimates when the tracking lost or degeneration occurs. Here we display the $z$, roll, and pitch components of poses since the $x$, $y$, and yaw components accurately fit with the ground truth. It can be observed that when tracking lost occurs (orange areas), the robot pose deviates significantly from the true value. Thanks to the proposed tracking recovery module, the pose can be immediately recovered to be close to the ground truth when streetlight observations arrive. The green area in Fig. 7 indicates that the robot experiences degenerated cases which result in pose drifting over time. When the robot observes streetlights without degeneration cases, the proposed strategy can effectively correct the drifted pose.
@@ V Conclusion
In this paper, a nocturnal vision-aided localization system integrating camera, IMU, and odometer sensors is proposed for robot navigation at night. To address problems of nocturnal visual localization, we propose to leverage streetlight detections for nocturnal localization. We utilize the InEKF for state estimation and propose novel data association and match extension modules to fully exploit the information from streetlights. Experiments on real nighttime scenes validate that the proposed system achieves accurate and robust localization which enables robots to autonomously navigate and explore in nighttime scenarios. Future work includes the integration with mainstream methods and other sensors for a more complete localization system which is adaptively available in both daytime and nighttime scenes.

### 03 Mar 2024 v2

@@ IV-A Implementation Details
The robot is equipped with a front-facing RGB camera ($30$ Hz $1280\times 720$ images), an odometer ($10$ Hz), and a Livox $360^{\circ}$ FOV LiDAR [^2] https://www.livoxtech.com/mid-360 ($10$ Hz) integrating an ICM40609 IMU ($200$ Hz). The camera’s intrinsics and all extrinsics among sensors have been pre-calibrated. All experiments were performed on the Intel i7-1165G7 CPU. Note that only an IMU, an odometer, and a monocular camera are used for the proposed nocturnal vision-aided localization system.
@@
Figure 5: Trajectories of ground truth (gray dashed line), VINS-Odom (green), and Night-Rider (blue) in four real scenes.
@@ IV-B Evaluation of Localization
To thoroughly evaluate the performance of the proposed system in nocturnal environments, we collect data in four different real scenes with multiple distribution types of streetlights. In the first scene (1149.2 m) and the fourth scene (723.8 m), the streetlights distribute sparsely on both sides of the road. Partial streetlights lie on only one side of the road in the second scene (1070.9 m). The third scene (777.2 m) contains densely distributed streetlights. All streetlight maps of these scenes are constructed using the method described in Section III-A. The offline optimized and precisely aligned trajectories from the loop closure augmented FAST-LIO2 using the LiDAR on board are treated as ground truths and EVO [28] is used for evaluation. Note that there are very few open-source research works targeting nocturnal scenarios. Thus we first compare Night-Rider with the state-of-the-art framework, extended VINS-Mono [3] with the odometer[29] (named VINS-Odom). As shown in Fig. 5, Night-Rider can precisely estimate trajectories with extremely low drifts in all four scenes while VINS-Odom drifts over time and even fails to track in the third scene. The reason is that the unstable illumination in nighttime scenes makes it difficult to extract reliable feature points from highly bright or dark images. This instability results in drifts or tracking failures of VINS-Odom. Contrarily, Night-Rider is independent of feature points and exploits the stable streetlights for localization. Thus trajectories of Night-Rider align accurately with ground truths. Table. I depicts the evaluation results of absolute trajectory error (ATE[30]) for translation and rotation parts. It shows that Night-Rider achieves remarkably accurate and robust localization in all scenes (<0.2% relative error of total trajectory length) and totally outperforms VINS-Odom, which further validates the effectiveness of our system.
@@
Figure 6: Trajectories in $x$, $y$ and $z$-axis of ground truth (gray dashed line), Night-Rider w/o Vision (red), Night-Rider w/o ME (green), and Night-Rider (blue) in four real scenes.
TABLE I: ATE[30] for translation and rotation of VINS-Odom, Night-Rider w/o Vision, Night-Rider w/o ME, and Night-Rider in four scenes. The best is shown in bold.
@@ IV-C Tracking Lost and Degeneration Cases
Fig. 7 shows the estimated trajectory of scene $3$ when the tracking lost or degeneration occurs. Here we display the $z$, roll, and pitch components of poses since the $x$, $y$, and yaw components accurately fit with the ground truth. The system without dealing with tracking lost or degeneration cases fail to estimate the robot poses. It can be observed that when tracking lost occurs (orange areas), the robot pose deviates significantly from the true value. Thanks to the proposed tracking recovery module, the pose can be immediately recovered to be close to the ground truth when streetlight observations arrive. The green area in Fig. 7 indicates that the robot experiences degenerated cases which result in pose drifting over time. When the robot observes streetlights without degeneration cases, the proposed strategy can effectively correct the drifted pose.
@@ V Conclusion
In this paper, a nocturnal vision-aided localization system integrating camera, IMU, and odometer sensors is proposed for robot navigation at night. To address problems of nocturnal visual localization, we propose to leverage streetlight detections for nocturnal localization. We utilize the InEKF for state estimation and propose novel data association and match extension modules to fully exploit the information from streetlights. Experiments on real nighttime scenes validate that the proposed system achieves accurate and robust localization. Future work includes the integration with mainstream methods and other sensors for a more complete localization system which is adaptively available in both daytime and nighttime scenes.
Acknowledgement
This work was supported in part by the Science and Technology Development Fund of Macao SAR under Grants 0046/2021/AGJ, 0067/2023/AFJ, 0123/2022/AFJ, and 0081/2022/A2.